OpenStudy (anonymous):

What is the remainder when \(\large 4444^{4444}\) is divided by \(\large 9\)?

OpenStudy (anonymous):

start taking digital remainders

OpenStudy (anonymous):

in the exponents \(4+4+4+4=16\) and \(1+6=7\) so the remainder when you divide \(4444\) by \(9\) is \(7\)

OpenStudy (anonymous):

therefore the remainder of \(4444^{4444}\) when divided by \(9\) is the same as the remainder of \(7^7\) when divided by \(9\)

OpenStudy (anonymous):

you have some more work to do to finish, but i hope that part was clear enough

OpenStudy (anonymous):

well not much actually \[7^7=49^3\times 7\] should finish it up nicely

OpenStudy (anonymous):

i see :) Thank you

ganeshie8 (ganeshie8):

nice :) \[\large 4444^{4444} \equiv (4\times 11\times 101)^{4444} \equiv 2^{4444} \mod 9\]